OpenStudy (anonymous):

One card is dealt. Find the chance that its color is red or its rank is an even number.

terenzreignz (terenzreignz):

Good catch, @Nurali I completely forgot about that :D

OpenStudy (anonymous):

Chance of an even number - 2,4,6,8,10 of 4 different colors chance of a red colour - 13*2=26 so 2,4,6,8 of 2 red colours are eliminated as they are contained in red colour cards so the probability is 26+ 10/52 that is 36/52

OpenStudy (perl):

P( red or even card ) = P(red ) + P(even) - P( red & even) = 26/52 + 20/52 - 10/52

OpenStudy (perl):

this is the addition rule P( A or B ) = P(A) + P(B) - P(A&B)